OpenStudy (anonymous):

Find the equation of the tangent line to the graph f(x)=3x-(24/x²) at the point where x=-2 Our teacher did not even give the correct answer soI don't know where I am at. Can anyone show me how I would go about answering this step by step??

zepdrix (zepdrix):

\[\Large f(x)\quad=\quad 3x-\frac{24}{x^2}\]

zepdrix (zepdrix):

We're looking for a straight line which is `tangent` to the our curve f(x) at x=-2. So it will be a line of the form:\[\Large y=mx+b\]Where m, the slope of the line, is given by the derivative of our function f(x) at x=-2.

zepdrix (zepdrix):

\[\Large m=f'(-2)\]

zepdrix (zepdrix):

Do you understand how to find \(\Large f'(x)\) ?

OpenStudy (anonymous):

Yes and no. I am still trying to work on the derivatives and I know I am missing small details when working on there type of problems

zepdrix (zepdrix):

Here's a handy rule of exponents.\[\Large \color{royalblue}{\frac{1}{x^n}\quad=\quad x^{-n}}\]

zepdrix (zepdrix):

Applying the rule to our function:\[\Large f(x)\quad=\quad 3x-24x^{-2}\]

zepdrix (zepdrix):

From this form, we can apply the power rule to each term.

zepdrix (zepdrix):

Power rule tells us to do two things: ~Bring the exponent down as a coefficient in front. ~Decrease the exponent by 1.

zepdrix (zepdrix):

\[\Large f(x)=3x^1-24x^{-2}\]So our derivative, applying the power rule to each term:\[\Large f'(x)=3(1\cdot x^0)-24(-2\cdot x^{-3})\]

zepdrix (zepdrix):

Understand how that works? The -2 comes down, then we subtract 1 from the exponent, decreasing it to -3.

OpenStudy (anonymous):

Yes I am with you so far!

zepdrix (zepdrix):

So this will simplify just a bit.\[\Large f'(x)=3+48x^{-3} \qquad\to\qquad f'(x)=3+\frac{48}{x^3}\]

zepdrix (zepdrix):

So the `slope` of our tangent line is given by the derivative function. So what do you get when you plug in x=-2?\[\Large f'(-2)=\;?\]

OpenStudy (anonymous):

15? I may be wrong I am just doing it in my head along with studying for my animal morphology test tomorrow lol

OpenStudy (anonymous):

sorry I leftout the negative

zepdrix (zepdrix):

Mmm I think it's -3ish

OpenStudy (anonymous):

Yes yes you are right. That negative I left out made a huge difference! ^^; Thanks so much you really helped me alot on understanding this!

zepdrix (zepdrix):

\[\Large m=f'(-2)=-3\] \[\Large y=mx+b \qquad\to\qquad y=-3x+\color{orangered}{b}\]

zepdrix (zepdrix):

So the find the equation of the line tangent to the curve, we need 2 things: ~The slope of the line ~The y-intercept So we've finished part 1.

zepdrix (zepdrix):

To find the y-intercept, we'll need to plug in a point that falls on the line.

zepdrix (zepdrix):

I've drawn an example above, hopefully this will help you understand the concept. The line tangent to the curve `touches` it at the one point.